# Interacting Crystalline Topological Insulators in 2D with Spin-Orbit Coupling

### v1

@@ II.3 Decorating Wyckoff patches
To find the classification of $p6mm$, we follow the procedure outlined in Sec. II: (i) we subdivide the unit cell into Wyckoff patches, (ii) we enumerate the decorations for all the Wyckoff patches, (iii) we then identify the gappable decorations, (iv) and we identify equivalent decorations.
@@
The unit cell subdivision is done based on the notion of Wyckoff patches. The Wyckoff positions and the site-symmetry group of the points in the unit cell are shown in Fig. 2(a),(b) and are listed in Tab. 4. The unit cell of this space group contains three 0D Wyckoff patches, $1a$, $2b$, and $3c$, two 1D Wyckoff patches, $6d$ and $6e$, and the 2D Wyckoff patch $12f$.
@@
The set of possible decorations of the $p6mm$ lattice alongside the indices that classify the decorations are summarized in Tab. 4 and depicted in Fig. 2(c)–(f). For the 0D Wyckoff patches, the indices describing the decorations are the total number of Kramers’ pairs at each Wyckoff patch
@@
$(n^{\,}_{6d},\,n^{\,}_{6e}\,m^{\,}_{6d},\,m^{\,}_{6e})\,\in(\mathbb{Z})^{2}\times(\mathbb{Z}^{\,}_{2})^{2}.$ (11)
@@
We now rule out those decorations that result in gapless modes appearing in the bulk, and we consider 0D, 1D and 2D decorations one by one. All the 0D decorations are gappable, therefore there is no reduction of any of the indices counting the Kramers’ pairs or mirror and rotation charges due to the requirement of gappability [Fig. 2(c)]. For 1D decorations on the $6e$ Wyckoff patches, there is only one minimal way to decorate the unit cells such that six lines meet at every crossing point, allowing to gap the edge modes in the bulk, while other decorations are not gappable: If we choose the $6e$ patches to terminate at the $2b$ patch, whose site-symmetry group contains mirror and three-fold rotation symmetries, there would be three 0D gapless boundary modes meeting at a point that are not gappable, and such a decoration has to be discarded [Fig. 2(e)]. However, if the $6d$ patches are continued to the next $1a$ position, the resulting decoration is gappable [Fig. 2(f)]. The 1D decorations along the $6d$ patches are symmetry constrained to meet in even number at every crossing point, therefore they can be gapped out in the bulk [Fig. 2(f)]. Likewise, the edge modes on neighboring 2D Wyckoff patches can be gapped out by covering the whole 2D space by the same phase [Fig. 2(d)].
@@
We now consider the equivalences that reduce the set of indices listed above by connecting different phases through adiabatic and symmetric transformations. We start by listing the processes of peeling off charges: For example, $6k$ Kramers’ pairs can be pulled away from the $1a$ Wyckoff patch, moved along the $6e$ lines, and brought to the $2b$ Wyckoff patches. This is illustrated in Fig. 2(g), (h). We write this process as
@@
From the collection of equivalences listed above, we identify the following independent indices
@@
and bearing in mind that there is still a well defined index counting the total number of Kramers’ pair in the unit cell, $n\in\mathbb{Z}$. Another set of reduction is realized by the neutralization of the mirror charges where an odd number of Wyckoff patches meet. In particular, the $m^{\,}_{2b}$ and $m^{\,}_{6e}$ indices reduce to trivial and therefore are dropped out in the classification [App. A].
@@
×$\mathbb{Z}\times\mathbb{Z}^{\,}_{6}\times\mathbb{Z}^{\,}_{3}\times(\mathbb{Z}^{\,}_{2})^{6}\times\mathbb{Z}^{\,}_{2}.$ (16)
The $(\mathbb{Z}^{\,}_{2})^{7}$ contribution is both intrinsically interacting and requires crystalline symmetries to exist.
All the remaining wallpaper groups are treated similarly, and the resulting classification is contained in Tab. 1, where entries corresponding to intrinsically interacting phases are highlighted. The key steps followed towards filling Tab. 1 are reported in App. B, where we also list the decoration indices corresponding to each wallpaper group classification explicitly.
@@ III Hamiltonian realizations
We present some of the possible Hamiltonian realizations for the 0D and 1D decorations. From the latter, one can construct toy Hamiltonians for the full 2D cFSPTs by taking the transitionally invariant sums of 0D and 1D Hamiltonians acting on the appropriate Wyckoff patches.
@@
Non-interacting electronic gapped and non-degenerate ground states with TRS and particle conservation are constrained to transform trivially under the action of crystalline symmetries [91]
@@
Beyond the example of the Hubbard square, this construction can be extended to other point groups and used to obtain cFSPTs as ground states of Hubbard Hamiltonians [98, 99]. The cFSPTs obtained as product states of 0D cFSPTs realized through Hubbard like Hamiltonians – so-called Mott Atomic Limits (MAL) – were studied in Ref. 99. There, the cFSPT wavefunctions are written as
@@
where each $\hat{O}^{\dagger}_{\bm{r}}$ contains $N$-particle creation operators, and $\hat{O}^{\dagger}_{\bm{r}}$’s commute at different sites, while they describe entangled states within the $N$-particles they create. Similar states are studied in Ref. 100.
@@
In Sec. II.2 we have pointed out that fermionic 1D systems with TRS, particle number conservation, and a mirror symmetry acting as an onSite-symmetry have $\mathbb{Z}^{\,}_{2}$ classification. In this section, we propose a model Hamiltonian that realizes the nontrivial phase of this classification, which we dub the AIIM model.
To better understand this $\mathbb{Z}^{\,}_{2}$ classification, we notice that the internal symmetry group in Eq. (4) (with $b=2$) has the following isomorphism
@@
with $\mathbb{Z}^{\mathrm{T}^{\prime}}_{2}$ being a spinless TRS whose generator is the product of the TRS and the mirror symmetry generators, $t\in\mathbb{Z}^{\mathrm{FT}}_{4}$ and $m\in\mathbb{Z}^{\mathrm{FM}}_{4}$, i.e., $t^{\prime}=t\,m$. Indeed, $t^{\prime}$ is represented by an antiunitary operator which squares to the identity, and has a nontrivial action on the particle conservation $\mathrm{U}(1)^{\mathrm{F}}$ subgroup. We claim that the 1D nontrivial FSPT phase with symmetry group (22) must be intrinsically interacting, meaning that there is no non-interacting counterpart of such a phase. The reason for this is as follows. The only nontrivial projective representation of group $G^{(1)}_{f,\mathrm{tot}}$ is when the representation $\widehat{U}(t^{\prime})$ of element $t^{\prime}$ squares to minus identity. In other words, the nontrivial 1D FSPT with $G^{(1)}_{f,\mathrm{tot}}$ symmetry must support gapless boundary modes that realize a projective representation of the subgroup
@@
which describes the symmetry class AI of the tenfold way. It is known that the non-interacting classification for class AI in 1D is trivial [1, 2, 3], while its fully interacting classification is $\mathbb{Z}^{\,}_{2}$ [6].
@@
where |$\ket{0}$ is the vacuum state. Hence, at low-energies the 16-dimensional local fermionic Fock space at each site $j$ is reduced to the 2-dimensional Hilbert space spanned by states |$\ket{1}^{\,}_{j}$ and |$\ket{2}^{\,}_{j}$. In this subspace, the effective Hamiltonian becomes
@@
which is just a complex phase. Fermion parity symmetry, $\widehat{U}(\theta=\pi)$, then just becomes identity, which is why the TRS and mirror symmetry subgroups $\mathbb{Z}^{\mathrm{FT}}_{4}$ and $\mathbb{Z}^{\mathrm{F}}_{4}$ reduce to $\mathbb{Z}^{\mathrm{T}}_{2}$ and $\mathbb{Z}^{\mathrm{M}}_{2}$, respectively. A gappable decoration of a 1D line by the AIIM Hamiltonian (25) then corresponds to the index $\nu=1$ introduced in Sec. II.2. While a trivial decoration corresponds to $\nu=0$.
@@
to the Hamiltonian, where operators with indices $j\leq L$ ($j\geq 1$) belong to the first (second) chain. If several pairs of 1D patches are related by rotation symmetry, as in the example of $6d$ and $6e$ decorations in $p6mm$ [Fig. 2(f)], a set of rotationally symmetric copies of (33) are added to the full Hamiltonian.
@@ IV Signatures of cFSPT phases
Figure 3: (a)-(b) Model Hamiltonians: (a) Schematic of the Hubbard square with the four-fold rotation center marked by a blue square, and the two hopping terms, and Hubbard interaction indicated. (b) Schematic of the AIIM chain. The action of a single term in the Hamiltonian in Eq. (25) is shown, alongside the mirror symmetry axis. In (a) and (b), the dots indicate spinful fermionic orbitals. (c)–(f) Distinguishing cFSPTs: (c) Square lattice ($p4mm$) decorated by the rotation-odd Hubbard square ground state of Eq. (20) at the $1b$ Wyckoff patch (the corner of the unit cell). Since there are no physical sites at the $1b$ Wyckoff patches, the lattice can be terminated by cutting through the decorations. This leads to a local two-fold degeneracy of the ground state at each corner, enforced by the combination of TRS, half filling, and $C_{4}$ symmetry of the full cluster. (d) Evaluation of the $C^{\,}_{4}$ charge at the $1a$ Wyckoff position in $p4mm$. (e) Example of string order parameter evaluation along the path $\gamma$ (marked by the pink line) on 1D Wyckoff patches’ decorations (marked by blue lines), which terminate with zero boundary modes for the nontrivial AIIM phase (marked by green dots). (f) A $\pi/2$ disclination in the $p4mm$ lattice leaves a gapless mode pinned at the disclination core.
@@ IV.1 Signatures for 0D decorations
As we have discussed in Sec. II.2 nontrivial 0D decorations are characterized by symmetry charges localized at 0D Wyckoff patches. For a given 2D lattice with $L$ unit cells in total, related by translations, the many-body ground state of the full lattice |$\ket{\Psi^{\,}_{L}}$ transforms as
@@
For cFSPTs that are only composed by 0D decorations, it was hinted in Refs. 104, 99 that single-particle Green’s function, often used to treat topological interacting systems by defining a ‘topological Hamiltonian’ [105, 106, 107, 108, 104], fail in distinguishing between trivial and nontrivial symmetry charges of 0D decorations. However, in Ref. 99 is was shown that 0D decorations where the nontrivial symmetry charge is carried by 2-particle entangled ‘clusters’ of electrons can be distinguished by the 2-particle Green’s function, and a generalization to $n$-particle entanglement and $n$-particle Green’s function is suggested. In Ref. 100 an alternative scheme based on real space invariants was proposed.
@@ IV.2 Signatures for 1D decorations
For cFSPTs with nontrivial 1D decorations, the ground state |$\ket{\Psi}$ does not necessarily carry global charge under symmetries. However, the 1D FSPT phases acquire an expectation value for non-local string order parameters [112, 113, 114]. Namely, the analogue of Eq. (34) is given by
@@
in terms of the effective spin-1/2 degrees of freedom introduced in Eq. (25c), where $j=1,\cdots,N$ labels the sites along the line $\gamma$. Alternatively, ground states of cFSPT phases can be detected by partial crystalline symmetry transformations [115, 116, 117].
There are two paradigmatic signatures of cFSPTs with nontrivial 1D decorations. First, when open boundary conditions are imposed, such cFSPTs feature gapless boundary states that are protected by both internal and crystalline symmetries. More precisely, as shown for the AIIM model in Sec. III.2, each 1D decoration features an effective spin-1/2 degree of freedom when open boundary conditions are imposed. The boundary is then described by TRS invariant spin-1/2 chains with translation and (if present) additional crystalline symmetries. The protected gaplessness at the boundary can then be understood as a direct consequence of the generalized Lieb-Schultz-Mattis (LSM) theorems [118, 119, 120, 121, 122, 123, 79, 124, 125, 126, 127, 128, 129] that apply to both effective spin-1/2 as well as the original fermionic degrees of freedom. Second, there are gapless boundary modes localized at lattice defects such as dislocations and disclinations (if rotation symmetry is present) [74]. This is because the otherwise gappable 1D decorations can no longer be generically gapped at the defect center as dislocations and disclinations require the removal of appropritate 1D decorations. For example, in the $p4mm$ wallpaper group, where four 1D Wyckoff patches meet at the center of the square, a $\pi/2$ disclination, which amounts to ‘cutting’ a $\pi/2$ angle wide slice from the lattice, would prevent the hybridization of one of the gapless spin-1/2 degrees of freedom, therefore leading to a gapless mode located at the disclination core [Fig. 3(f)]. We refer to the mode pinned at the defect core as gapless, in the sense that it leads to a true ground state degeneracy, arising from local degrees of freedom.

### v2

@@ II.3 Decorating Wyckoff patches
To find the classification of $p6mm$, we follow the procedure outlined in Sec. II: (i) we subdivide the unit-cell into Wyckoff patches, (ii) we enumerate the decorations for all the Wyckoff patches, (iii) we then identify the gappable decorations, (iv) and we identify equivalent decorations.
@@
The unit-cell subdivision is done based on the notion of Wyckoff patches. The Wyckoff positions and the site-symmetry group of the points in the unit-cell are shown in Fig. 2(a),(b) and are listed in Table 4. The unit-cell of this space group contains three 0D Wyckoff patches, $1a$, $2b$, and $3c$, two 1D Wyckoff patches, $6d$ and $6e$, and the 2D Wyckoff patch $12f$.
@@
The set of possible decorations of the $p6mm$ lattice alongside the indices that classify the decorations are summarized in Table 4 and depicted in Figs. 2(c)–(f). For the 0D Wyckoff patches, the indices describing the decorations are the total number of Kramers’ pairs at each Wyckoff patch
@@
$(n^{\,}_{6d},\,n^{\,}_{6e},\,m^{\,}_{6d},\,m^{\,}_{6e})\,\in(\mathbb{Z})^{2}\times(\mathbb{Z}^{\,}_{2})^{2}.$ (11)
@@
We now rule out those decorations that result in gapless modes appearing in the bulk, and we consider 0D, 1D and 2D decorations one by one. All the 0D decorations are gappable, therefore there is no reduction of any of the indices counting the Kramers’ pairs or mirror and rotation charges due to the requirement of gappability [Fig. 2(c)]. For 1D decorations on the $6e$ Wyckoff patches, there is only one minimal way to decorate the unit-cells such that six lines meet at every crossing point, allowing to gap the edge modes in the bulk, while other decorations are not gappable: If we choose the $6e$ patches to terminate at the $2b$ patch, whose site-symmetry group contains mirror and three-fold rotation symmetries, there would be three 0D gapless boundary modes meeting at a point that are not gappable, and such a decoration has to be discarded [Fig. 2(e)]. However, if the $6d$ patches are continued to the next $1a$ position, the resulting decoration is gappable [Fig. 2(f)]. The 1D decorations along the $6d$ patches are symmetry constrained to meet in even number at every crossing point, therefore, they can be gapped out in the bulk [Fig. 2(f)]. Likewise, the edge modes on neighboring 2D Wyckoff patches can be gapped out by covering the whole 2D space by the same phase [Fig. 2(d)].
@@
We now consider the equivalences that reduce the set of indices listed above by connecting different phases through adiabatic and symmetric transformations. We start by listing the processes of peeling off charges: For example, $6k$ Kramers’ pairs can be pulled away from the $1a$ Wyckoff patch, moved along the $6e$ lines, and brought to the $2b$ Wyckoff patches. This is illustrated in Figs. 2(g) and 2(h). We write this process as
@@
From the collection of equivalences listed above, we identify the following independent indices:
@@
and bearing in mind that there is still a well defined index counting the total number of Kramers’ pair in the unit-cell, $n\in\mathbb{Z}$. Another set of reduction is realized by the neutralization of the mirror charges where an odd number of Wyckoff patches meet. In particular, the $m^{\,}_{2b}$ and $m^{\,}_{6e}$ indices reduce to trivial and therefore are dropped out in the classification (Appendix A).
@@
×$\begin{split}\mathbb{Z}\times\mathbb{Z}^{\,}_{6}\times\mathbb{Z}^{\,}_{3}\times(\mathbb{Z}^{\,}_{2})^{6}\times\mathbb{Z}^{\,}_{2}.\end{split}$ (16)
In this classification, the $(\mathbb{Z}^{\,}_{2})^{6}$ contribution, with indices $(r_{1a},\,m_{1a},\,r_{3c},\,m_{3c},\,\nu_{6d},\,\nu_{6e})\in(\mathbb{Z}^{\,}_{2})^{6}$, is both intrinsically interacting and requires crystalline symmetries to exist. For this reason, they are highlighted in cyan in Table 1.
All the remaining wallpaper groups are treated similarly, and the resulting classification is contained in Table 1, where entries corresponding to intrinsically interacting phases are highlighted. The key steps followed towards filling Table 1 are reported in Appendix B, where we also list the decoration indices corresponding to each wallpaper group classification explicitly.
@@ III Hamiltonian realizations
We present some of the possible Hamiltonian realizations for the 0D and 1D decorations. From the latter, one can construct toy Hamiltonians for the full 2D cFSPTs by taking the transitionally invariant sums of 0D and 1D Hamiltonians acting on the appropriate Wyckoff patches. We do not discuss the case of 2D Wyckoff patches, as there are no crystalline symmetries acting onsite, and the $\mathbb{Z}_{2}$ classification that arises without crystalline symmetries. An example of realization of the nontrivial phase is for instance the Kane-Mele model, for the case of noninteracting fermions [95].
@@
Noninteracting electronic gapped and non-degenerate ground states with TRS and particle conservation are constrained to transform trivially under the action of crystalline symmetries [91]
@@
Beyond the example of the Hubbard square, this construction can be extended to other point groups and used to obtain cFSPTs as ground states of Hubbard Hamiltonians [99, 100]. The cFSPTs obtained as product states of 0D cFSPTs realized through Hubbard like Hamiltonians, so-called Mott atomic limits (MAL), were studied in Ref. 100. There, the cFSPT wave functions are written as
@@
where each $\hat{O}^{\dagger}_{\bm{r}}$ contains $N$-particle creation operators, and $\hat{O}^{\dagger}_{\bm{r}}$’s commute at different sites, while they describe entangled states within the $N$-particles they create. Similar states are studied in Ref. 101.
@@
In Sec. II.2 we have pointed out that fermionic 1D systems with TRS, particle-number conservation, and a mirror symmetry acting as an onsite symmetry have $\mathbb{Z}^{\,}_{2}$ classification. In this section, we propose a model Hamiltonian that realizes the nontrivial phase of this classification, which we dub the AIIM model.
To better understand this $\mathbb{Z}^{\,}_{2}$ classification, we notice that the internal symmetry group in Eq. (4) (with $b=2$) has the following isomorphism:
@@
with $\mathbb{Z}^{\mathrm{T}^{\prime}}_{2}$ being a spinless TRS whose generator is the product of the TRS and the mirror symmetry generators, $t\in\mathbb{Z}^{\mathrm{FT}}_{4}$ and $m\in\mathbb{Z}^{\mathrm{FM}}_{4}$, i.e., $t^{\prime}=t\,m$. Indeed, $t^{\prime}$ is represented by an antiunitary operator which squares to the identity, and has a nontrivial action on the particle conservation $\mathrm{U}(1)^{\mathrm{F}}$ subgroup. We claim that the 1D nontrivial FSPT phase with symmetry group (22) must be intrinsically interacting, meaning that there is no noninteracting counterpart of such a phase. The reason for this is as follows. The only nontrivial projective representation of group $G^{(1)}_{f,\mathrm{tot}}$ is when the representation $\widehat{U}(t^{\prime})$ of element $t^{\prime}$ squares to minus identity. In other words, the nontrivial 1D FSPT with $G^{(1)}_{f,\mathrm{tot}}$ symmetry must support gapless boundary modes that realize a projective representation of the subgroup
@@
which describes the symmetry class AI of the 10-fold way. It is known that the noninteracting classification for class AI in 1D is trivial [1, 2, 3], while its fully interacting classification is $\mathbb{Z}^{\,}_{2}$ [6].
@@
where |$\ket{0}$ is the vacuum state. Hence, at low-energies the 16-dimensional local fermionic Fock space at each site $j$ is reduced to the 2D Hilbert space spanned by states |$\ket{1}^{\,}_{j}$ and |$\ket{2}^{\,}_{j}$. In this subspace, the effective Hamiltonian becomes
@@
which is just a complex phase. Fermion parity symmetry, $\widehat{U}(\theta=\pi)$, then just becomes identity, which is why the TRS and mirror symmetry subgroups $\mathbb{Z}^{\mathrm{FT}}_{4}$ and $\mathbb{Z}^{\mathrm{F}}_{4}$ reduce to $\mathbb{Z}^{\mathrm{T}}_{2}$ and $\mathbb{Z}^{\mathrm{M}}_{2}$, respectively. A gappable decoration of a 1D line by the AIIM Hamiltonian (25) then corresponds to the index $\nu=1$ introduced in Sec. II.2, while a trivial decoration corresponds to $\nu=0$.
@@
to the Hamiltonian. In Eq. (33), operators with indices $j\leq L$ belong to the first chain, and with $j\geq 1$ to the second, and they are distinguished by non calligraphic and calligraphic symbols, respectively. If several pairs of 1D patches are related by rotation symmetry, as in the example of $6d$ and $6e$ decorations in $p6mm$ [Fig. 2(f)], a set of rotationally symmetric copies of (33) are added to the full Hamiltonian.
@@ IV Signatures of cFSPT phases
Figure 3: (a), (b) Model Hamiltonians: (a) schematic of the Hubbard square with the four-fold rotation center marked by a blue square, and the two hopping terms, and Hubbard interaction indicated. (b) Schematic of the AIIM chain. The action of a single term in the Hamiltonian in Eq. (25) is shown, alongside the mirror-symmetry axis. In (a) and (b), the dots indicate spinful fermionic orbitals. (c)–(f) Distinguishing cFSPTs: (c) Square lattice ($p4mm$) decorated by the rotation-odd Hubbard square ground state of Eq. (20) at the $1b$ Wyckoff patch (the corner of the unit-cell). Since there are no physical sites at the $1b$ Wyckoff patches, the lattice can be terminated by cutting through the decorations. This leads to a local two-fold degeneracy of the ground state at each corner, enforced by the combination of TRS, half-filling, and $C_{4}$ symmetry of the full cluster. (d) Evaluation of the $C^{\,}_{4}$ charge at the $1a$ Wyckoff position in $p4mm$. (e) Example of string order-parameter evaluation along the path $\gamma$ (marked by the pink line) on 1D Wyckoff patches’ decorations (marked by blue lines), which terminate with zero boundary modes for the nontrivial AIIM phase (marked by green dots). (f) A $\pi/2$ disclination in the $p4mm$ lattice leaves a gapless mode pinned at the disclination core.
@@ IV.1 Signatures for 0D decorations
As we have discussed in Sec. II.2 nontrivial 0D decorations are characterized by symmetry charges localized at 0D Wyckoff patches. For a given 2D lattice with $L$ unit-cells in total, related by translations, the many-body ground state of the full lattice |$\ket{\Psi^{\,}_{L}}$ transforms as
@@
For cFSPTs that are only composed by 0D decorations, it was hinted in Refs. 105, 100 that single-particle Green’s function, often used to treat topological interacting systems by defining a “topological Hamiltonian” [106, 107, 108, 109, 105], fail in distinguishing between trivial and nontrivial symmetry charges of 0D decorations. However, in Ref. 100 is was shown that 0D decorations where the nontrivial symmetry charge is carried by two-particle entangled “clusters” of electrons can be distinguished by the two-particle Green’s function, and a generalization to $n$-particle entanglement and $n$-particle Green’s function is suggested. In Ref. 101 an alternative scheme based on real space invariants was proposed.
@@ IV.2 Signatures for 1D decorations
For cFSPTs with nontrivial 1D decorations, the ground state |$\ket{\Psi}$ does not necessarily carry global charge under symmetries. However, the 1D FSPT phases acquire an expectation value for non-local string order parameters [113, 114, 115]. Namely, the analog of Eq. (34) is given by
@@
in terms of the effective spin-$\frac{1}{2}$ degrees of freedom introduced in Eq. (25c), where $j=1,\cdots,N$ labels the sites along the line $\gamma$. Alternatively, ground states of cFSPT phases can be detected by partial crystalline symmetry transformations [116, 117, 118].
There are two paradigmatic signatures of cFSPTs with nontrivial 1D decorations. First, when open boundary conditions are imposed, such cFSPTs feature gapless boundary states that are protected by both internal and crystalline symmetries. More precisely, as shown for the AIIM model in Sec. III.2, each 1D decoration features an effective spin-1/2 degree of freedom when open boundary conditions are imposed. The boundary is then described by TRS invariant spin-$\frac{1}{2}$ chains with translation and (if present) additional crystalline symmetries. The protected gaplessness at the boundary can then be understood as a direct consequence of the generalized Lieb-Schultz-Mattis (LSM) theorems [119, 120, 121, 122, 123, 124, 79, 125, 126, 127, 128, 129, 130] that apply to both effective spin-1/2 as well as the original fermionic degrees of freedom. Second, there are gapless boundary modes localized at lattice defects such as dislocations and disclinations (if rotation symmetry is present) [74]. This is because the otherwise gappable 1D decorations can no longer be generically gapped at the defect center as dislocations and disclinations require the removal of appropriate 1D decorations. For example, in the $p4mm$ wallpaper group, where four 1D Wyckoff patches meet at the center of the square, a $\pi/2$ disclination, which amounts to “cutting” a $\pi/2$ angle wide slice from the lattice, would prevent the hybridization of one of the gapless spin-1/2 degrees of freedom, therefore leading to a gapless mode located at the disclination core [Fig. 3(f)]. We refer to the mode pinned at the defect core as gapless, in the sense that it leads to a true ground state degeneracy, arising from local degrees of freedom.